the market value of CCC's equity is $15 million and the market value is of its debt it $5 million. If the required rates of return on the equity is 12% and that on its debt is 8%, calculate the company's cost of capital. CCC's tax rate is 20%
Answer #1


solution: WACC=rd*(1-t)*wd+re*we WACC=0.08*(1-0.2)* (5/(5+15))+0.12*(15/(5+15)) WACC= 10.60% rd Pretax cost of debt wd weight